# package/before/amsmath

### 14 Dec 2023 v1

@@ 1 Introduction
The pre-commitment solution to the mean-variance problem often reformulates the original mean-variance problem to an equivalent formulation Li and Ng [2000]; Forsyth [2011]; Bjork et al. [2017]. Pre-commitment strategy works well within stable economic regimes and is shown to outperform time-consistent strategies Forsyth [2020]; Vigna [2020]. However, the pre-commitment assumes that there is complete information from $t$ to $T$ but in reality the state of the economy might change drastically within a given time period due to exogenous factors. The absence of time consistency also prevents us from using dynamic programming directly.
@@
To overcome the time inconsistency problem of the mean variance criteria, there has been increasing development in continuous time consistent mean variance asset allocation in both reinsurance and investments Guan and Hu [2022]. The current approaches typically solve the continuous time optimal portfolio allocation problem Zhou and Li [2000]; Aivaliotis and Veretennikov [2018]; Aivaliotis and Palczewski [2014]; Wang and Forsyth [2011]. The approach in Van Staden et al. [2018] considers the control of impulse controls in jump processes.
@@
In discrete time, this problem has been solved using dynamic programming assuming local optimality Basak and Chabakauri [2010]. The discrete time solution does not solve for the value function of the Hamilton-Jacobi-Bellman (HJB) equation. The analytical solution to the continuous time consistent mean variance asset allocation problem has been studied for an agent with $d$-assets Zhou and Li [2000] and in the $1$ asset reinsurance and investment setting using $K$ agents Guan and Hu [2022]. Note, the optimal asset allocation problem can be viewed as a special case of the optimal trade execution problem when there is no price impact, i.e. $\kappa_{t}=0$ and $\kappa_{P}=0$ for agents who can only sell/buy the asset. The numerical solutions to the optimal trade execution problem has been presented for the time-inconsistent strategy using finite difference methods (FDM) Forsyth [2011]; Wang and Forsyth [2011]. The FDM solution to the time-consistent mean variance asset allocation problem, i.e. control on $\alpha$, and its extension to impulse control has also been solved Van Staden et al. [2018].
@@
Recent advances in reinforcement learning (RL) uses model free methods to solve the optimal trade execution on limit-order-books and real data Ning et al. [2021]; Lin and Beling [2021]; Chen et al. [2022]. A RL framework that solves the continuous time mean-variance portfolio optimization problem derives the optimal policy iteration Wang and Zhou [2020]. However, this only considers the control of the portfolio allocation and not the optimal trade rate. A RL method has been used to approximate the optimal trade execution under mean variance criteria in the discrete time setting Hendricks and Wilcox [2014]. Typically, RL methods use the model free (Q-learning) approach. However, the issue with using Q-learning in the optimal trade execution in discrete time is that the problem is solved locally and does not guarantee the solution given satisfies the HJB equation Basak and Chabakauri [2010]. In current frameworks, there is no clear environment where agents can interact with. The agents learn independently based on a general simulated market. In this paper, we propose a multi-agent model which couples each agent $k$ implicitly through the price of the asset. This creates an environment that takes into account the decisions made by each individual agent and allows us to simulate the market dynamics more precisely.
@@
    We formulate the mean variance optimal trade execution problem such that the investor is aware of other investors in the market. We reformulate the mean variance problem into its dual representation and an auxillary HJB equation. To extend this framework into higher dimensions, we reformulate the HJB equation as a backward stochastic differential equation (BSDE). The auxillary HJB equation has been shown to have a viscosity solution Aivaliotis and Veretennikov [2018].
@@
    We present a neural network solution for the $d$-dimensional assets and $K$-agent optimal trade execution problem using neural networks, which extends the numerical solution of the time consistent optimal trade execution problem to high dimensions. We use techniques based on neural network pricing of high dimensional options to formulate a deep residual self-attention network that overcomes the need to save weights at each timestep E et al. [2017]; Chen and Wan [2021]; Han and E [2016]; Hure et al. [2020].
@@ 2 Mean-Variance Optimal Trade Execution
In this section, we present the optimal trade execution problem with $i=1,...,d$ assets and $K$-agents. This is an extension of the classical framework Forsyth [2011]. Suppose each agent $k$, where $k\in[1,...,K]$, can observe a basket of $d$ stocks in the market; with price process $\underline{S}_{k}(t)=[S_{1,k}(t),...,S_{d,k}(t)]^{\top}\in\mathbb{R}^{d}$. Then the number of shares of the $d$ stocks owned by agent $k$ is given by $\underline{\alpha}_{k}(t)=[\alpha_{1,k}(t),...,\alpha_{d,k}(t)]$ and the amount invested in a risk free bank account is given by $b_{k}\in\mathbb{R}$. Let $t\in[0,T]$ be the time up to maturity $T$, then at any time $t$, the market agent $k$ has a wealth $X_{k}(t)$ given by:
@@
To handle buying and selling cases symmetrically, we start off with $b_{0}$ in cash, $\alpha_{0}>0$ shares if we are selling and $\alpha_{0}<0$ shares if buying. This means that we are trying to liquidate a long (short) position if we are selling (buying). More precisely
@@
where $b_{T}$ is the cash generated by selling/buying in $[0,T)$. The objective is to maximize the wealth at maturity, $T$, and simultaneously minimize the risk as measured by the variance. Note that due to the non-monotonicity of the variance term, this problem as presented is time inconsistent Basak and Chabakauri [2010]; Tse [2012].
A natural optimization criteria to consider for optimal trade execution is the mean variance (MV) strategy of Markowitz. The MV strategy looks for the optimal strategy that maximizes an investors expected gain, and at the same time minimizes their risk Wang and Forsyth [2011].
In this section we define the mean-variance criteria for optimal trade execution. The agent $k$ with trade rate $v_{k}(t)$ maximizes the expected wealth defined as $\mathbb{E}_{v}[X_{k}(T)]=\mathbb{E}[X_{k}(T);v_{k}(t)]$ and minimizes the variance defined as
$Var_{v}[X_{k}(T)]=\mathbb{E}_{v}[X(T)^{2}]-(\mathbb{E}_{v}[X_{k}(T)])^{2}.$
@@
$\textnormal{arg}\max\limits_{v}J(X_{k}(T))=\textnormal{arg}\max\limits_{v}\left\{\mathbb{E}_{v}[X_{k}(T)]-\frac{\gamma}{2}Var_{v}(X_{k}(T))\right\}.$ (5)
@@
$J^{*}(X_{k}(T))=\mathbb{E}_{v^{*}}[X_{k}(T)]-\frac{\gamma}{2}Var_{v^{*}}(X_{k}(T)).$
In general, the solution for the mean-variance problem is time inconsistent. Many approaches have been explored to overcome this limitation by solving an auxiliary problem which is referred to as the time-consistent mean variance problem Zhou and Li [2000]; Forsyth [2011]; Tse [2012]; Aivaliotis and Palczewski [2014]; Van Staden et al. [2018]; Forsyth [2020]; Guan and Hu [2022].
@@ 2.1 Problem Formulation
The dynamics of the optimal trade execution problem can be formulated by the Almgren-Chriss model Almgren and Chriss [2000]. The instantaneous trade rate is given by:
@@
where $\alpha(t)$ is the amount of stock held by a liquidating agent. Let $\mu$ be the drift, $r$ be the risk free rate and $\sigma$ be the volatility of the asset. The price process of the risky asset follows the Geometric Brownian motion (GBM) Forsyth [2011]:
@@
where $g(v(t))$ is the level of permanent price impact. The cash position is then given by Forsyth [2011]:
$\frac{db(t)}{dt}=rb(t)-v(t)s(t)f(v(t)),$
where $f(v(t))$ is the level of temporary price impact. We can define the temporary price change due to the temporary impact as $S_{temp}(t,v)=v(t)s(t)f(v(t))$, $S_{temp}(v,t)\leq S(t)$.
Given states $\{S(T^{-}),B(T^{-}),\alpha(T^{-})\}$ at the instant $t=T^{-}$ before the end of the trading horizon, we have one final liquidation (if $\alpha(T^{-})\neq 0$) to ensure that the number of shares $\alpha(T)=0$. The liquidation value is computed as follows Tse [2012]:
$B(T)$ $=B(T^{-})+\lim_{v\rightarrow-\infty}\{\alpha(T^{-})S_{temp}(T^{-},v)\}$
$=\int_{t=0}^{T^{-}}-e^{r(T-t)}v(t)S_{temp}(t,v)dt+\lim_{v\rightarrow-\infty}\{\alpha(T^{-})S_{temp}(T^{-},v)\}.$
@@ 2.2 Extension of the Almgren-Chriss model to $d$ Assets and $K$ agents
In this section, we describe how we extend the dynamics of the optimal trade execution problem to $d$ assets and $K$ agents. First, we consider $d$ assets and one agent. Let the trade rate for each asset $i$ be given by:
$v_{i}(t)=\frac{d\alpha_{i}(t)}{dt},$ (6)
we can define the the aggregate trade rate for an agent as $v(t)=\sum_{i}^{d}v_{i}(t)$. As with the one dimensional case, we assume the process $S_{i}(t)$ follows the GBM, with a modification due to the permanent price impact. Let $\mu_{i}$ and $\sigma_{i}$ be the drift and volatility of stocks $i=1,...,d$. Let $\mathbf{\rho}\in\mathbb{R}^{d\times d}$ be the correlation matrix and we define the correlated random variable $dW_{i}(t)=\sum\limits_{j=1}^{d}L_{i,j}Z_{j}\sqrt{dt}$, where $Z_{j}\sim N(0,1)$ are i.i.d, and $\mathbf{L}$ is the Cholesky factor of $\mathbf{\rho}$, i.e. $\mathbf{\rho}=\mathbf{LL}^{\top}$.
@@
For simplicity, we assume that the market has sufficient liquidity and we do not consider market frictions. A multi-asset optimal trade execution problem has been modelled by a Orstein-Uhlenbeck (OU) process Bergault et al. [2022]. In our model, we assume a constant $\mu_{i}$, $r$ and $\sigma_{i}$ and we introduce a permanent price impact function $g(v_{i}(t))$ in (7). We use the following linear form for the permanent price impact
$g(v_{i}(t))=\kappa_{p}v_{i}(t),$
where the constant $\kappa_{p}$ is the permanent price impact factor. A linear permanent price impact function is used to eliminate round-trip arbitrage opportunities Tse [2012].
Given a stock price $\underline{s}=[s_{1},...,s_{d}]^{\top}$, The bank account $b$ is assumed to follow
$\frac{db(t)}{dt}=rb(t)-v(t)\sum\limits_{i=1}^{d}S_{i}(t)f(v_{i}(t)),$ (8)
where $f(v_{i}(t))$ is the temporary price impact of the investor as they sell/buy the asset $i$. The function $f(v_{i})$ is assumed to have the form
$f(v_{i})=[1+\kappa_{s,i}\textnormal{sign}(v_{i})]\textnormal{exp}\{\kappa_{\tau}\textnormal{sign}(v_{i})|v_{i}|^{\beta}\},$
where $\kappa_{s,i}$ is the bid-ask spread for each asset $i$, $\kappa_{\tau}$ is the temporary price impact factor of the investor and $\beta$ is the price impact exponent Forsyth [2011].
Next, we extend the optimal trade execution problem to $K$-agents. Let $k=1,...,K$, then the trading rate for each agent $k$ and each asset $i$ is given by
$v_{i,k}(t)=\frac{d\alpha_{i,k}(t)}{dt},$ (9)
and the aggregate trade rate for each agent $k$ is given by:
$v_{k}(t)=\sum_{i=1}^{d}v_{i,k}(t).$
Formally, we introduce the trade impacts in the price of asset $i$ as
$dS_{i}(t)=(\mu_{i}-r+\sum_{k}g(v_{i,k}(t)))S_{i}(t)dt+\sigma_{i}S_{i}(t)dW_{i}(t),$
where $g(v_{i,k}(t))=\kappa_{p,k}\textnormal{sign}(v_{i,k}(t))|v_{i,k}(t)|$. This couples the actions of each agent $k$ implicitly through the price dynamics. Note, since our economy or state space is modeled through the asset prices, it is a way for each agent’s actions to update the state space. This also means the economy has full information about each agent $k$. In this formulation, the trade rate of each trader $k$ affects other traders through the permanent price impact on each asset $i$. This tells us that the asset price is driven by not just the drift of our asset but also by the accumulation of permanent price impact of each agent $k$. This is consistent with the observation that in a multiple agent setting the total market impact on an asset $i$ is the sum of permanent price impact of all agents and the temporary price impact Schoneborn [2008].
@@
$f(v_{i,k}(t))=[1+\kappa_{s,i}\textnormal{sign}(v_{i,k}(t))]\textnormal{exp}\{\kappa_{\tau,k}\textnormal{sign}(v_{i,k}(t))|v_{i,k}(t)|^{\beta_{k}}\}.$
The temporary price impact factors of each agent $k$ only affects the execution price of agent $k$, thus there is no coupling involved in this term with other agents. This gives us a measure of how each agent perceives their own impact on the market.

### 19 Mar 2024 v2

@@ 1 Introduction
The pre-commitment solution to the mean-variance problem often reformulates the original mean-variance problem to an equivalent formulation [Li and Ng 2000; Forsyth 2011; Bjork et al. 2017]. Pre-commitment strategy works well within stable economic regimes and is shown to outperform time-consistent strategies [Forsyth 2020; Vigna 2020]. However, the pre-commitment assumes that there is complete information from $t$ to $T$ but in reality the state of the economy might change drastically within a given time period due to exogenous factors. The absence of time consistency also prevents us from using dynamic programming directly.
@@
To overcome the time inconsistency problem of the mean variance criteria, there has been increasing development in continuous time consistent mean variance asset allocation in both reinsurance and investments [Guan and Hu 2022]. The current approaches typically solve the continuous time optimal portfolio allocation problem [Zhou and Li 2000; Aivaliotis and Veretennikov 2018; Aivaliotis and Palczewski 2014; Wang and Forsyth 2011]. The approach in [Van Staden et al. 2018] considers the control of impulse controls in jump processes.
@@
In discrete time, this problem has been solved using dynamic programming assuming local optimality [Basak and Chabakauri 2010]. The discrete time solution does not solve for the value function of the Hamilton-Jacobi-Bellman (HJB) equation. The analytical solution to the continuous time consistent mean variance asset allocation problem has been studied for an agent with $d$-assets [Zhou and Li 2000] and in the $1$ asset reinsurance and investment setting using $K$ agents [Guan and Hu 2022]. Note, the optimal asset allocation problem can be viewed as a special case of the optimal trade execution problem when there is no price impact, i.e. $\kappa_{t}=0$ and $\kappa_{P}=0$ for agents who can only sell/buy the asset. The numerical solutions to the optimal trade execution problem has been presented for the time-inconsistent strategy using finite difference methods (FDM) [Forsyth 2011; Wang and Forsyth 2011]. The FDM solution to the time-consistent mean variance asset allocation problem, i.e. control on $\alpha$, and its extension to impulse control has also been solved [Van Staden et al. 2018].
@@
Recent advances in reinforcement learning (RL) uses model free methods to solve the optimal trade execution on limit-order-books and real data [Ning et al. 2021; Lin and Beling 2021; Chen et al. 2022]. A RL framework that solves the continuous time mean-variance portfolio optimization problem derives the optimal policy iteration [Wang and Zhou 2020]. However, this only considers the control of the portfolio allocation and not the optimal trade rate. A RL method has been used to approximate the optimal trade execution under mean variance criteria in the discrete time setting [Hendricks and Wilcox 2014]. Typically, RL methods use the model free (Q-learning) approach. However, the issue with using Q-learning in the optimal trade execution in discrete time is that the problem is solved locally and does not guarantee the solution given satisfies the HJB equation [Basak and Chabakauri 2010]. In current frameworks, there is no clear environment where agents can interact with. The agents learn independently based on a general simulated market. In this paper, we propose a multi-agent model which couples each agent $k$ implicitly through the price of the asset. This creates an environment that takes into account the decisions made by each individual agent and allows us to simulate the market dynamics more precisely.
@@
    We formulate the mean variance optimal trade execution problem such that the investor is aware of other investors in the market. We reformulate the mean variance problem into its dual representation and an auxillary HJB equation. To extend this framework into higher dimensions, we reformulate the HJB equation as a backward stochastic differential equation (BSDE). The auxillary HJB equation has been shown to have a viscosity solution [Aivaliotis and Veretennikov 2018].
@@
    We present a neural network solution for the $d$-dimensional assets and $K$-agent optimal trade execution problem using neural networks, which extends the numerical solution of the time consistent optimal trade execution problem to high dimensions. We use techniques based on neural network pricing of high dimensional options to formulate a deep residual self-attention network that overcomes the need to save weights at each timestep [E et al. 2017; Chen and Wan 2021; Han and E 2016; Hure et al. 2020].
@@ 2 Mean-Variance Optimal Trade Execution
In this section, we present the optimal trade execution problem with $i=1,...,d$ assets and $K$-agents. This is an extension of the classical framework [Forsyth 2011]. Suppose each agent $k$, where $k\in[1,...,K]$, can observe a basket of $d$ stocks in the market; with price process $\underline{S}_{k}(t)=[S_{1,k}(t),...,S_{d,k}(t)]^{\top}\in\mathbb{R}^{d}$. Then the number of shares of the $d$ stocks owned by agent $k$ is given by $\underline{\alpha}_{k}(t)=[\alpha_{1,k}(t),...,\alpha_{d,k}(t)]$ and the amount invested in a risk free bank account is given by $b_{k}\in\mathbb{R}$. Let $t\in[0,T]$ be the time up to maturity $T$, then at any time $t$, the market agent $k$ has a wealth $X_{k}(t)$ given by:
@@
To handle buying and selling cases symmetrically, we start off with $b_{0}\in\mathbb{R}$ in cash, $\alpha_{0}>0$ shares if we are selling and $\alpha_{0}<0$ shares if buying. This means that we are trying to liquidate a long (short) position if we are selling (buying). More precisely
@@
where $b_{T}-b_{0}$ is the cash generated by selling/buying in $[0,T)$. The objective is to maximize the wealth at maturity, $T$, and simultaneously minimize the risk as measured by the variance. Note that due to the non-monotonicity of the variance term, this problem as presented is time inconsistent [Basak and Chabakauri 2010; Tse 2012].
A natural optimization criteria to consider for optimal trade execution is the mean-variance (MV) strategy of Markowitz. The MV strategy looks for the optimal strategy that maximizes an investors expected gain, and at the same time minimizes their risk [Wang and Forsyth 2011].
In this section we define the mean-variance criteria for optimal trade execution. Let $\mathbb{E}[X_{k}(T)|v_{k}(t)]$ be the conditional expected wealth given the trade rate $v_{k}(t)$. We denote $\mathbb{E}_{v}[X_{k}(T)]=\mathbb{E}[X_{k}(T)|v_{k}(t)]$. The agent $k$ with trade rate $v_{k}(t)$ maximizes the expected wealth and minimizes the variance defined as
$Var_{v}[X_{k}(T)]=\mathbb{E}_{v}[X_{k}(T)^{2}]-(\mathbb{E}_{v}[X_{k}(T)])^{2}.$
@@
$\textnormal{arg}\max\limits_{v}J(X_{k}(T))=\textnormal{arg}\max\limits_{v}\left\{\mathbb{E}_{v}[X_{k}(T)]-\frac{\gamma_{k}}{2}Var_{v}(X_{k}(T))\right\}.$ (5)
@@
$J^{*}(X_{k}(T))=\mathbb{E}_{v^{*}}[X_{k}(T)]-\frac{\gamma_{k}}{2}Var_{v^{*}}(X_{k}(T)).$
In general, the solution for the mean-variance problem is time inconsistent. Many approaches have been explored to overcome this limitation by solving an auxiliary problem which is referred to as the time consistent mean-variance problem [Zhou and Li 2000; Forsyth 2011; Tse 2012; Aivaliotis and Palczewski 2014; Van Staden et al. 2018; Forsyth 2020; Guan and Hu 2022].
@@ 2.1 Problem Formulation
The dynamics of the optimal trade execution problem can be formulated by the Almgren-Chriss model [Almgren and Chriss 2000]. The instantaneous trade rate is given by:
@@
where $\alpha(t)$ is the amount of stock held by a liquidating agent. Let $\mu$ be the drift, $r$ be the risk free rate and $\sigma$ be the volatility of the asset. The price process of the risky asset follows the Geometric Brownian motion (GBM) [Forsyth 2011]:
@@
where $g(v(t))$ is the level of permanent price impact. We use the following linear form for the permanent price impact [gatheral-2010; Forsyth 2011]:
$g(v(t))=\kappa_{p}v(t),$
where the constant $\kappa_{p}$ is a constant permanent price impact factor.
The stock price we receive from trading is given by the price we actually receive from a trade is given by [Forsyth 2011]:
$\bar{S}(t,v(t))=S(t)f(v(t)),$
where $f(v(t))$ is the level of temporary price impact. The temporary price impact is used in our model to capture liquidity effects on the trade [almgren-2012]. The function $f(v(t))$ is assumed to have the form
$f(v(t))=[1+\kappa_{s}\textnormal{sign}(v(t))]\textnormal{exp}\{\kappa_{\tau}v(t)^{\beta}\},$
where $\kappa_{s}$ is the bid-ask spread for the asset, $\kappa_{\tau}$ is a constant that represents the temporary price impact factor of the investor and $\beta$ is the price impact exponent [Forsyth 2011].
The cash position is then given by [Forsyth 2011]:
$\frac{db(t)}{dt}=rb(t)-v(t)S(t)f(v(t)).$
Given states $\{S(T^{-}),B(T^{-}),\alpha(T^{-})\}$ at the instant $t=T^{-}$ before the end of the trading horizon, we have one final liquidation (if $\alpha(T^{-})\neq 0$) to ensure that the number of shares $\alpha(T)=0$. The liquidation value is computed as follows [Tse 2012]:
$B(T)$ $=B(T^{-})+\lim_{v\rightarrow-\infty}\{\alpha(T^{-})\bar{S}(T^{-},v(T^{-}))\}$
$=\int_{t=0}^{T^{-}}-e^{r(T-t)}v(t)\bar{S}(t,v(t))dt+\lim_{v\rightarrow-\infty}\{\alpha(T^{-})\bar{S}(T^{-},v(T^{-})\}.$
@@ 2.2 Extension of the Almgren-Chriss model to $d$ Assets and $K$ agents
In this section, we describe how we extend the dynamics of the optimal trade execution problem to $d$ assets and $K$ agents. First, we consider $d$ assets and one agent. Let the trade rate for the agent with $d$ assets be given by $\underline{v}(t)=[v_{1}(t),...,v_{d}(t)]^{\top}$. As with the one dimensional case, we assume the process $S_{i}(t)$ follows the GBM, with a modification due to the permanent price impact. Let $\mu_{i}$ and $\sigma_{i}$ be the drift and volatility of stocks $i=1,...,d$. Let $\mathbf{\rho}\in\mathbb{R}^{d\times d}$ be the correlation matrix and we define the correlated random variable $dW_{i}(t)=\sum\limits_{j=1}^{d}L_{i,j}Z_{j}\sqrt{dt}$, where $Z_{j}\sim N(0,1)$ are i.i.d, and $\mathbf{L}$ is the Cholesky factor of $\mathbf{\rho}$, i.e. $\mathbf{\rho}=\mathbf{LL}^{\top}$.
@@
For simplicity, we assume that the market has sufficient liquidity and we do not consider market frictions. A multi-asset optimal trade execution problem has been modelled by a Orstein-Uhlenbeck (OU) process [Bergault et al. 2022]. In our model, we assume a constant $\mu_{i}$, $r$ and $\sigma_{i}$ and we introduce a permanent price impact function $g_{i}(v_{i}(t))$ in (6). We use the following linear form for the permanent price impact [Forsyth 2011]:
$g_{i}(v_{i}(t))=\kappa_{p,i}v_{i}(t),$
where the constant $\kappa_{p,i}$ is the permanent price impact factor for asset $i$. A linear permanent price impact function is used to eliminate round-trip arbitrage opportunities as shown in [Tse 2012]. Given a vector of stock prices $\underline{s}=[s_{1},...,s_{d}]^{\top}$, the bank account $b$ is assumed to follow
$\frac{db(t)}{dt}=rb(t)-\sum\limits_{i=1}^{d}v_{i}(t)S_{i}(t)f_{i}(v_{i}(t)),$ (7)
where $f_{i}(v_{i}(t))$ is the temporary price impact of the investor as they sell/buy the asset $i$. The function $f_{i}(v_{i}(t))$ is assumed to have the form
$f_{i}(v_{i}(t))=[1+\kappa_{s,i}\textnormal{sign}(v_{i}(t))]\textnormal{exp}\{\kappa_{\tau,i}v_{i}(t)^{\beta}\},$
where $\kappa_{s,i}$ is the bid-ask spread for each asset, $\kappa_{\tau,i}$ is the temporary price impact factor of the investor for each asset and $\beta$ is the price impact exponent [Forsyth 2011].
Next, we extend the optimal trade execution problem to $K$-agents. Let $k=1,...,K$, then the trading rate for each agent $k$ is given by $\underline{v}_{k}=[v_{1,k},...,v_{d,k}]^{\top}$. Let $S_{i}(t)$ be the price of asset $i$ that follows the GBM. For multiple agents the dynamics of asset $i$ is given by:
$dS_{i}(t)=(\mu_{i}-r+\sum_{k=1}^{K}g_{i,k}(v_{i,k}(t)))S_{i}(t)dt+\sigma_{i}S_{i}(t)dW_{i}(t),$
where $g_{i,k}(v_{i,k})=\kappa_{p,i,k}v_{i,k}$. The permanent price impact over multiple agents impact the price of asset $i$ as they liquidate their position in the asset. This couples the actions of each agent $k$ implicitly through the price dynamics. Note, since our economy or state space is modeled through asset prices. This also means the economy has full information about each agent $k$. This tells us that the asset price is driven by not just the drift of our asset but also by the trade activity of each agent $k$. This is consistent with the observation that in a multiple agent setting the total market impact on an asset $i$ is the sum of permanent price impact of all agents and the temporary price impact [Schoneborn 2008].
@@
$f_{i,k}(v_{i,k}(t))=[1+\kappa_{s,i}\textnormal{sign}(v_{i,k}(t))]\textnormal{exp}\{\kappa_{\tau,i,k}v_{i,k}(t)^{\beta_{k}}\}.$
Note, the bid-ask spread $\kappa_{s,i}$ is only dependent on the asset $i$. The temporary price impact factors of each agent k only affects the execution price of agent $k$, thus there is no coupling involved in this term with other agents. This gives us a measure of how each agent perceives their own impact on the market.